# DeepEye-SQL: A Software-Engineering-Inspired Text-to-SQL Framework

### v2

@@ 5.2. Overall Performance
OmniSQL (Li et al., 2025a) Qwen2.5-Coder-32B-Instruct $\sim$32B 67.0 89.8
@@
OpenSearch-SQL (v2) (Xie et al., 2025) GPT-4o >200B 69.3 87.1
Alpha-SQL (Li et al., 2025b) Qwen2.5-Coder-32B-Instruct $\sim$32B 69.7 -
@@
To answer this, we present the main performance of DeepEye-SQL on the BIRD-Dev and Spider-Test datasets in Table 2, comparing it against a wide range of state-of-the-art fine-tuning and prompting-based methods. The results clearly demonstrate that our software-grounded framework establishes a new benchmark for training-free Text-to-SQL generation.
Specifically, when integrated with the Qwen3-Coder-30B-A3B model, DeepEye-SQL achieves an execution accuracy of 73.5% on BIRD-Dev. This result not only significantly outperforms all existing prompting-based baselines but, more remarkably, it also surpasses the leading fine-tuning-based systems like XiYan-SQL (73.3%) and CHASE-SQL (73.0%). It is crucial to note that these competing methods rely on substantially larger and often proprietary models (e.g., GPT-4o and Gemini-1.5-Pro). On the Spider-Test dataset, DeepEye-SQL with Qwen3-Coder-30B-A3B achieves an execution accuracy of 89.8%. This performance matches the best fine-tuned method, OmniSQL, and surpasses other strong competitors like XiYan-SQL (89.7%). This demonstrates that our prompting-based framework can attain the same level of accuracy as highly specialized, fine-tuned models on this foundational benchmark without incurring any training costs.
Furthermore, the consistently high performance of DeepEye-SQL across all three tested open-source models—71.1% with Gemma3-27B, 70.9% with Qwen2.5-Coder-32B, and 73.5% with Qwen3-Coder-30B on BIRD, which underscores the robustness and generalizability of our framework.
@@
Table 3. Performance comparison with same LLMs.
Methods EX (%) Delta (%)
Gemma3-27B-Instruction
CoT-Baseline 54.2 -
CHESS (IR,CG,UT) 66.1 +11.9
Alpha-SQL 68.6 +14.4
DeepEye-SQL (Ours) 71.1 +16.9
Qwen2.5-Coder-32B-Instruct
CoT-Baseline 61.3 -
CHESS (IR,CG,UT) 67.7 +6.4
Alpha-SQL 69.7 +8.4
DeepEye-SQL (Ours) 70.9 +9.6
Qwen3-Coder-30B-A3B-Instruct
CoT-Baseline 61.7 -
CHESS (IR,CG,UT) 67.9 +6.2
Alpha-SQL 71.2 +9.5
DeepEye-SQL (Ours) 73.5 +11.8
@@ 5.3. Robust Schema Linking Analysis
Table 4. Schema linking analysis with Qwen3-Coder-30B-A3B on BIRD-Dev dataset.
Schema Linking Methods
Table
Recall (%)
Column
Recall (%)
# Avg.
Tokens
No Schema Linking - - 5486.2
@@ 5.4. Analysis on N-Version Programming for SQL Generation
Table 5. SQL generation analysis with Qwen3-Coder-30B-A3B on BIRD-Dev dataset.
SQL Generation EX (%) UB-EX (%)
@@
Figure 5. Correctness overlap analysis of three SQL generation methods using Qwen3-Coder-30B-A3B model on BIRD-Dev dataset.:
@@ 5.5. Confidence-aware SQL Selection Analysis
Figure 6. Execution accuracy vs. SQL confidence on BIRD-Dev dataset with Qwen3-Coder-30B-A3B model.:
Table 6. Performance comparison with different SQL selection methods on BIRD-Dev dataset.
Selection Methods BIRD-EX (%) Spider-EX (%)
@@
Confidence-aware SQL Selection 71.1 ($\uparrow$1.0) 88.9 ($\uparrow$0.6)
@@
Confidence-aware SQL Selection 70.9 ($\uparrow$0.7) 88.7 ($\uparrow$0.5)
@@
Confidence-aware SQL Selection 73.5 ($\uparrow$0.8) 89.8 ($\uparrow$0.1)
Figure 7. Execution accuracy vs. Confidence Shorcut Threshold on BIRD-Dev dataset with Qwen3-Coder-30B-A3B model.:
RQ5: Can a selection mechanism that leverages voting confidence outperform traditional consistency-based voting, and what is the motivation for such a design?
To answer this, we first analyze the relationship between voting confidence and correctness, and then present a direct comparison of our method against the standard self-consistency approach.
The core motivation behind our Confidence-aware SQL Selection is the observation that the reliability of consistency-based voting is highly dependent on the degree of consensus among the generated candidates. To validate this premise, we visualized the SQL confidence score (defined as the proportion of candidates in the largest execution cluster) against the actual execution accuracy for our results on the BIRD-Dev dataset. As visualized in Figure 6, there is a remarkably strong positive correlation between these two factors, with a Pearson correlation coefficient (Sedgwick, 2012) of 0.954. This result confirms that when the confidence score is high, the top-ranked SQL is very likely to be correct. Conversely, in low-confidence scenarios where candidates produce many different results, standard voting is unreliable. This insight directly informs our two-path design: a “shortcut” for high-confidence cases and a more discerning “full review” for low-confidence ones.
To quantify the effectiveness of this design, we compared its performance against traditional Consistency-based Voting (Li et al., 2025b; Gao et al., 2024) across three different base models. The results are shown in Table 6. For every model and on both the BIRD and Spider datasets, our method yields a notable improvement in execution accuracy. For instance, with Gemma3-27B-Instruct, our method improves the EX on BIRD by 1.0%, from 70.1% to 71.1%. This analysis validates that by identifying and re-evaluating ambiguous, low-confidence scenarios instead of blindly trusting the majority vote, our method achieves a more robust and accurate final selection.
RQ6: How sensitive is the framework’s performance to the choice of the confidence shortcut threshold ($\theta_{conf}$)?
To evaluate the robustness of our selection mechanism, we conducted a sensitivity analysis on the confidence shortcut threshold, $\theta_{conf}$. This threshold determines the trade-off between taking the efficient shortcut and triggering the full, LLM-based review. As shown in Figure 7, we varied $\theta_{conf}$ from 0.0 (always perform full review) to 1.0 (always take the shortcut) and observed the impact on execution accuracy. The results demonstrate that our framework’s performance is robust across a wide range of threshold values. The accuracy remains high and stable, peaking at 73.5% with a threshold of 0.6, and staying above 73.0% for all values between 0.4 and 1.0. This indicates that our method is not overly sensitive to the precise choice of this hyperparameter.
@@ 5.6. Ablation Study
Table 7. Ablation study of DeepEye-SQL with Qwen3-Coder-30B-A3B on BIRD-Dev dataset.
Configuration EX (%) Delta (%)
DeepEye-SQL 73.5 -
@@
- w/o Skeleton-based SQL Generation 72.2 -1.3
- w/o ICL-based SQL Generation 71.0 -2.5
- w/o D&C-based SQL Generation 72.3 -1.2
- w/o SQL Unit Testing and Revision via Tool-Chain 71.4 -2.1
- w/o Confidence-aware SQL Selection 72.7 -0.8
RQ7: What is the contribution of each key component to the overall performance of the DeepEye-SQL framework?
@@ 5.7. Efficiency Analysis
Table 8. Efficiency Comparison on BIRD-Dev with Qwen3-Coder-30B-A3B model.
Methods
Avg. Input
Tokens (K)
Avg. Output
Tokens (K)
EX (%)
CHESS (IR,SS,CS) 327.02 27.83 67.9
Alpha-SQL 138.03 72.21 71.2
DeepEye-SQL 23.21 23.16 73.5
- Semantic Value Retrieval 0.67 0.03 -
- Robust Schema Linking 13.91 6.33 -
- N-version Programming for SQL Generation 5.28 11.38 -
- SQL Unit Testing and Revision via Tool-Chain 3.16 5.41 -
- Confidence-aware SQL Selection 0.19 0.01 -
RQ8: How does DeepEye-SQL’s efficiency, in terms of token consumption, compare to other state-of-the-art prompting-based frameworks?
To answer this, we conducted a cost study to evaluate the token efficiency of our framework. While system latency is an important practical concern, it is highly sensitive to external factors such as hardware and deployment configuration. Therefore, we use token efficiency (both input and output) as a more objective and reproducible metric for the cost of LLM-based systems.
@@ 6. Conclusion
In this paper, we presented DeepEye-SQL, which reframes Text-to-SQL as a verifiable SDLC-style workflow. By unifying semantic grounding, N-version programming, deterministic tool-chain verification, and confidence-aware selection, DeepEye-SQL enforces end-to-end correctness and achieves system-level reliability. Extensive experiments validate that, using $\sim$30B open-source LLMs without fine-tuning, DeepEye-SQL achieves 73.5% EX on BIRD-Dev and 89.8% on Spider-Test, outperforming the state of the art. Our results suggest that principled orchestration offers a promising path toward system-level reliability in Text-to-SQL, beyond relying on LLM scaling alone.

### v3

@@ 4.2. Overall Performance
OmniSQL (Li et al., 2025b) Qwen2.5-Coder-32B-Instruct $\sim$32B 67.0 89.8
@@
OpenSearch-SQL (v2) (Xie et al., 2025a) GPT-4o >200B 69.3 87.1
Alpha-SQL (Li et al., 2025a) Qwen2.5-Coder-32B-Instruct $\sim$32B 69.7 -
@@
Table 4. Execution Accuracy (%) on Spider 2.0 Benchmark. All methods utilize DeepSeek-R1 as the backend LLM to ensure a fair comparison.
Methods Spider 2.0-Snow Spider 2.0-Lite
DIN-SQL (Pourreza and Rafiei, 2023) 0.0 4.8
DAIL-SQL (Gao et al., 2024) 6.6 8.8
Spider-Agent (Lei et al., 2025) 10.6 13.7
ReFoRCE (Deng et al., 2025) 38.0 29.6
RSL-SQL (Cao et al., 2024) - 30.5
LinkAlign (Wang and Liu, 2025) - 33.1
DeepEye-SQL (Ours) 50.5 38.2
@@
To answer this, we present the main performance of DeepEye-SQL on the BIRD-Dev, Spider-Test, and Spider 2.0 datasets in Table 3 and Table 4, comparing it against a wide range of state-of-the-art fine-tuning, prompting-based, and agentic methods. The results clearly demonstrate that our software-grounded framework establishes a new benchmark for Text-to-SQL generation. Specifically, when integrated with the Qwen3-Coder-30B-A3B model, DeepEye-SQL achieves an execution accuracy of 73.5% on BIRD-Dev. This result not only significantly outperforms all existing prompting-based baselines but, more remarkably, it also surpasses the leading fine-tuning-based systems like XiYan-SQL (73.3%) and CHASE-SQL (73.0%). It is crucial to note that these competing methods rely on substantially larger and often proprietary models (e.g., GPT-4o and Gemini-1.5-Pro). On the official BIRD held-out Test set, DeepEye-SQL achieves 75.07% execution accuracy with open-source models. Furthermore, scaling up to Gemini-3-Pro yields 76.58% (Global Rank 5th) even with minimal sampling ($N=1$). On the Spider-Test dataset, DeepEye-SQL with Qwen3-Coder-30B-A3B achieves an execution accuracy of 89.8%. This performance matches the best fine-tuned method, OmniSQL, and surpasses other strong competitors like XiYan-SQL (89.7%). This demonstrates that our prompting-based framework can attain the same level of accuracy as highly specialized, fine-tuned models on this foundational benchmark without incurring any training costs. Moreover, on the newly released enterprise-grade Spider 2.0 benchmark (Table 4), DeepEye-SQL demonstrates superior robustness in handling complex, real-world constraints. On the rigorous “Snow” subset, DeepEye-SQL achieves 50.5% execution accuracy, surpassing the specialized agentic framework ReFoRCE (38.0%) by a large margin (+12.5%). Notably, traditional prompting methods like DIN-SQL fail completely on this task (0.0%), highlighting that our structured engineering lifecycle is essential for navigating the ambiguity of realistic data warehousing scenarios. Furthermore, the consistently high performance of DeepEye-SQL across all three tested open-source models—71.1% with Gemma3-27B, 70.9% with Qwen2.5-Coder-32B, and 73.5% with Qwen3-Coder-30B on BIRD, which underscores the robustness and generalizability of our framework (Table 5).
@@
Table 5. Performance comparison using the same LLM backbone. Gemma3-27B, Qwen2.5-32B, and Qwen3-30B denote Gemma3-27B-Instruction, Qwen2.5-Coder-32B-Instruct, and Qwen3-Coder-30B-A3B-Instruct, respectively.
Methods Gemma3-27B Qwen2.5-32B Qwen3-30B
EX $\Delta$ EX $\Delta$ EX $\Delta$
CoT-Baseline 54.2 61.3 61.7
CHESS (IR, CG, UT) 66.1 +11.9 67.7 +6.4 67.9 +6.2
Alpha-SQL 68.6 +14.4 69.7 +8.4 71.2 +9.5
DeepEye-SQL (Ours) 71.1 +16.9 70.9 +9.6 73.5 +11.8
@@ 4.3. Robust Schema Linking Analysis
Table 6. Schema linking analysis on the BIRD-Dev dataset.
Schema Linking Methods Table Column Avg.
Recall (%) Recall (%) Tokens
No Schema Linking 5486.2
@@ 4.4. Analysis on N-Version Programming for SQL Generation
Table 7. Analysis of SQL generation methods on the BIRD-Dev dataset.
SQL Generation Method EX (%) UB-EX (%)
@@
Figure 5. Correctness overlap analysis of three SQL generation methods on BIRD-Dev dataset.
@@
RQ5: Is the N-version programming robust to variations in few-shot example quality?
Table 8. Sensitivity analysis of example quality with Qwen3-Coder-30B-A3B on the BIRD-Dev dataset.
Generator Retrieval Strategy SQL Jaccard EX (%) UB-EX (%)
Single ICL-based Least Relevant 0.44 70.1 75.8
Most Relevant 0.58 70.9 78.3
Hybrid (ICL+D&C+Skl) Least Relevant 0.44 71.5 80.2
Most Relevant 0.58 71.7 81.1
DeepEye-SQL (Full Pipeline) Least Relevant 0.44 73.4 80.8
Most Relevant 0.58 73.5 81.6
To answer this, we explicitly tested the system’s sensitivity by manipulating the retrieval strategy. Instead of the standard approach that selects the top-$K$ examples with the highest similarity (“Most Relevant”), we deliberately retrieved the $K$ examples with the lowest similarity (“Least Relevant”). We quantified the quality of these examples using the Jaccard similarity between the example SQL and the ground truth (following DAIL-SQL (Gao et al., 2024)). As shown in Table 8, the “Least Relevant” set indeed exhibits significantly lower quality (0.44 vs. 0.58). Under this challenging condition, the performance of the standalone ICL generator drops noticeably (70.9% $\to$ 70.1%). However, our Hybrid Generator dampens this drop (71.7% vs. 71.5%), and the full DeepEye-SQL demonstrates remarkable stability (73.5% vs. 73.4%), verifying that our SE-inspired orchestration effectively masks the volatility of retrieval quality.
@@ 4.5. Confidence-aware SQL Selection Analysis
Figure 6. Execution accuracy vs. SQL confidence on BIRD-Dev dataset with Qwen3-Coder-30B-A3B model.
Table 9. Performance comparison of different SQL selection methods on the BIRD-Dev and Spider datasets.
Selection Method BIRD-EX (%) Spider-EX (%)
@@
Confidence-aware selection 71.1 ($\uparrow$1.0) 88.9 ($\uparrow$0.6)
@@
Confidence-aware selection 70.9 ($\uparrow$0.7) 88.7 ($\uparrow$0.5)
@@
Confidence-aware selection 73.5 ($\uparrow$0.8) 89.8 ($\uparrow$0.1)
Figure 7. Execution accuracy vs. Confidence Shorcut Threshold on BIRD-Dev dataset with Qwen3-Coder-30B-A3B model.
RQ6: Effectiveness of Confidence-aware Selection.
We first validate that voting reliability strongly correlates with consensus. As shown in Figure 6, we observe a high Pearson correlation of 0.954 between voting confidence and execution accuracy on BIRD-Dev. This confirms that while high-confidence votes are reliable (justifying our efficient “shortcut”), low-confidence scenarios require the deeper re-evaluation our method provides. Comparative results in Table 9 demonstrate that our approach consistently outperforms traditional Consistency-based Voting (Li et al., 2025a; Gao et al., 2024). For instance, with Gemma3-27B, we achieve a 1.0% gain (70.1% $\to$ 71.1%) on BIRD, proving that selectively reviewing ambiguous cases yields superior robustness over blind majority voting.
RQ7: Sensitivity to Confidence Threshold ($\theta_{conf}$).
We evaluated the robustness of the threshold $\theta_{conf}$, which balances efficiency against rigorous review. Figure 7 shows that performance remains highly stable across a wide range: accuracy stays above 73.0% for $\theta_{conf}\in[0.4,1.0]$, peaking at 73.5% when $\theta_{conf}=0.6$. This indicates that our framework is not sensitive to specific hyperparameter choices.
@@ 4.6. Ablation Study
Table 10. Ablation study of DeepEye-SQL with Qwen3-Coder-30B-A3B on the BIRD-Dev dataset.
Variant EX (%) $\Delta$ (%)
DeepEye-SQL 73.5
@@
w/o Skeleton-based Generation 72.2 -1.3
w/o ICL-based Generation 71.0 -2.5
w/o D&C-based Generation 72.3 -1.2
w/o Tool-Chain Testing & Revision 71.4 -2.1
w/o Confidence-aware Selection 72.7 -0.8
RQ8: What is the contribution of each key component to the overall performance of the DeepEye-SQL framework?
@@ 4.7. Efficiency Analysis
Table 11. Efficiency comparison on the BIRD-Dev dataset. DeepEye-SQL achieves orders-of-magnitude lower latency than search-based (Alpha-SQL) and linear-chain (CHESS) baselines while maintaining higher accuracy.
Method Avg. Input Avg. Output Avg. Latency EX (%)
Tokens (K) Tokens (K) (s/query)
CHESS (IR, SS, CS) 327.02 27.83 284.4 67.9
Alpha-SQL 138.03 72.21 377.1 71.2
DeepEye-SQL 23.21 23.16 17.8 73.5
Semantic Value Retrieval 0.67 0.03
Robust Schema Linking 13.91 6.33
N-version Generation 5.28 11.38
Tool-Chain Verification 3.16 5.41
Confidence-aware Selection 0.19 0.01
@@
RQ9: How does DeepEye-SQL perform in terms of efficiency and scalability compared to SOTA baselines?
To answer this, we conducted a comprehensive evaluation covering token cost, end-to-end latency, and system scalability.
[…]
(iii) Scalability Analysis. To further assess robustness, Figure 8 breaks down performance by database complexity.
  • Robust Accuracy (Figure 8 Left): DeepEye-SQL maintains high execution accuracy across Small, Medium, and Large schemas, showing no significant degradation even as complexity increases.
[…]
Figure 9. Comparative Error Analysis on BIRD-Dev.
4.8. Error Analysis
RQ10: What are the dominant failure modes of DeepEye-SQL, and how does its error distribution compare to SOTA baselines?
To gain deeper insights into the limitations of DeepEye-SQL, we performed a manual inspection of failure cases on the BIRD-Dev set. We compared DeepEye-SQL with OmniSQL (Li et al., 2025b) and Alpha-SQL (Li et al., 2025a) using a fine-grained taxonomy covering five error dimensions, each divided into specific subclasses: Schema (① Selection, ② Join), Value (① Mapping, ② Format), Function (① Aggregation, ② Calculation), Logic (① Filter, ② Presentation), and Ground-truth (① Gold Error, ② Ambiguity).
Figure 9 illustrates the comparative distribution of these errors. DeepEye-SQL demonstrates a substantial advantage in handling Value-related errors (only 17 instances), significantly outperforming baselines (56 and 39). This attributes to our Semantic Value Retrieval strategy which effectively resolves both ① Value Mapping and ② Data Format issues. Furthermore, the lower error rate in Logic and Function categories (67 and 75 respectively) validates the synergy of our N-version SQL Generation and SQL Unit Testing & Revision modules. However, Schema Linking remains the most challenging category for all models (126 errors for DeepEye-SQL). Failures here are dominated by incorrect ① Column Selection and ② Join Paths, typically triggered by complex multi-hop requirements or obscure column names. We also observed that roughly 30% of the reported failures (122 cases) fall into the Ground-truth category, stemming from ambiguity in questions or errors in the gold SQLs, indicating the ceiling imposed by dataset quality.
5. Related Work
Text-to-SQL Solutions. The task of translating natural language into executable SQL queries has evolved significantly (Liu et al., 2025a). The recent emergence of Large Language Models (LLMs) has marked a new paradigm. Current research on applying LLMs to Text-to-SQL is largely divided into two categories: fine-tuning and prompting-based methods. Fine-tuning methods adapt open-source models like Code Llama or Qwen for the Text-to-SQL task by training them on large corpora of question-SQL pairs (Li et al., 2024b; Yang et al., 2024; Li et al., 2025b; Sheng et al., 2025; Liu et al., 2025c). This approach can yield highly efficient and specialized models but may exhibit limited generalization to out-of-domain or highly complex scenarios. In contrast, prompting-based methods leverage the powerful in-context learning and reasoning capabilities of very large, often proprietary, models like GPT-4o (Hurst et al., 2024) and Gemini (Anil et al., 2023) without requiring model training. To manage the complexity of the task, these methods typically decompose the process into a multi-stage pipeline, including sub-tasks like schema linking, SQL generation, and refinement (Pourreza and Rafiei, 2023; Li et al., 2024a; Pourreza et al., 2025; Li et al., 2025a; Ma et al., 2024; Zhu et al., 2025a). While powerful, these frameworks often rely on a single generation path and the fallible self-correction abilities of the LLM, which can compromise robustness. Our work, DeepEye-SQL, belongs to the prompting-based category but differentiates itself by explicitly adopting a systematic framework inspired by software engineering.
Software Engineering. The challenge of building robust and reliable LLM-based systems often mirrors the complexities of traditional software development. In response, software engineering has established a set of core principles to manage complexity and ensure product quality. The Software Development Life Cycle (SDLC) (Ruparelia, 2010) provides a foundational, systematic process for software creation, typically involving phases such as requirements analysis, implementation, testing, and deployment. To deconstruct existing systems and inform new designs, practitioners often employ Reverse Engineering (Rekoff, 1985), a process of analyzing a finished product to deduce its underlying specifications. To enhance system reliability, fault tolerance techniques are critical. A notable example is N-Version Programming (Chen and Avizienis, 1978), where multiple, independently developed versions of a component are executed in parallel, and their results are adjudicated to mask faults and increase the likelihood of a correct outcome. For quality assurance, Unit Testing (Runeson, 2006) is a fundamental practice. It involves the granular testing of individual software components or “units” in isolation to verify that each part functions correctly according to its design specifications. This bottom-up approach is crucial for identifying and rectifying defects early in the development process. Finally, to govern the release process, a Quality Gate (Schnell and Reinhart, 2016) acts as a final checkpoint, enforcing a set of predefined criteria to determine whether a software artifact meets the required quality standard for deployment. Although central to traditional software engineering, these principles have seen limited systematic adoption in LLM-based pipelines. We bring them to Text-to-SQL to enable a structured, verifiable, and robust pipeline.
@@ 6. Conclusion
In this paper, we present DeepEye-SQL, a verifiable SDLC-style workflow for Text-to-SQL that targets system-level reliability through end-to-end correctness. Without any fine-tuning, DeepEye-SQL leverages open-source MoE LLMs (about 30B total parameters with roughly 3B activated) and achieves 73.5% execution accuracy on BIRD-Dev, 75.07% on the official BIRD-Test leaderboard, and 89.8% on Spider-Test, outperforming state-of-the-art approaches that rely on larger models or extensive training. Overall, these results suggest that principled orchestration, rather than fine-tuning or scaling up inference alone, is a promising path toward reliable Text-to-SQL.
Acknowledgements.
This paper was supported by the NSF of China (62402409); Youth S&T Talent Support Programme of Guangdong Provincial Association for Science and Technology (SKXRC2025461); the Young Talent Support Project of Guangzhou Association for Science and Technology (QT-2025-001); Guangdong Basic and Applied Basic Research Foundation (2023A1515110545); Guangzhou Basic and Applied Basic Research Foundation (2025A04J3935); and Guangzhou-HKUST(GZ) Joint Funding Program (2025A03J3714).